The expression 1/3x^2-2 can be rewritten as 1/3(x-k)(x+k),where k is a positive constant.What is the value of k?
Answer by stanbon(75887)

The expression 1/3x^2-2 can be rewritten as 1/3(x-k)(x+k),where k is a positive constant.What is the value of k?
-----
(1/3)x^2-2 = (1/3)(x^2 - 6) = (1/3)(x-sqrt(6)(x+sqrt(6))
------------------------------------
k = sqrt(6)
--------------
Cheers,
Stan H.
-------------